My daughter is helping her neice on an algebra problem. It has been a while since she has done algebra. She loves to help tutor students but is stumped.
How do you work the problem:
3x - y = 3

Answer by elima(1433)

Begin by making a table choosing values for x then solving the equation for y;
x|y
----
1|
2|
0|
When x=1;
3(1)-y=3
3-y=3
y=3-3
y=0
When x=2
3(2)-y=3
6-y=3
-y=3-6
-y=-3
y=3
When x=0;
3(0)-y=3
0-y=3
-y=3
y=-3
Hope you understand
=)